OpenStudy (anonymous):

how would you solve 4cos2A = 3cosA for 90º≤A≤180º.

OpenStudy (thomas5267):

\[ 4\cos(2A)=4\cos^2(A)-4\sin^2(A)=4\cos^2(x)-4\left(1-\cos^2(x)\right)\\ =8\cos^2(x)-4 \]